OpenStudy (anonymous):

A bag contains 5 white marbles,4 black marble and 1 red marble.Find the probability of picking a red marble? 1.1/10 2.1/2 3.2/5 4.1/5

OpenStudy (anonymous):

Total marbles = 5+4+1 = 10 How many red marbles? 1. P(red marble) = (# of red marbles)/(# of total marbles) P(red) = 1/10 Your answer is #1.

OpenStudy (anonymous):

Thanks